The length of a rectangle exceed its width by 6 units. If each dimension is increased by 3 units, the area would be increased by 57 square units. Find the dimension of the rectangle
Answer by richwmiller(17219)

l=w+6,
a=l*w,
(l+3)*(w+3)=a+57,
b=a+57
a = 55, b = 112,
original rectangle L = 11, W = 5